# DeepVoting: Learning Voting Rules with Tailored Embeddings

### v1

@@ 7 Conclusions and Future Work
In this paper we have shown that not only can we efficiently learn known PSCFs from preference data, but also that we can modify these rules in order to improve them in ways that, to date, have not be possible through traditional algorithmic design methods. We have highlighted the importance of the choice of embedding on the efficiency and quality of the learned rules, finding some surprising results including that rules like IRV, Borda, and Black’s Rule, which are not majoritarian rules, can be learned well from $T_{WT}$. It remains to be seen whether other embeddings can be designed, of size $m\times m$ or smaller, that outperform the embeddings we took from the social choice literature. Different embeddings may be beneficial in particular for rules whose outcomes are NP-Hard to compute or other common axioms.
@@ Acknowledgments
This work was funded in part by GoodLynx and IBM. This material is based upon work supported by the National Science Foundation under Grant #2030859 to the Computing Research Association for the CIFellows Project. Ben Abramowitz was supported by the CIFellows Project and GoodLynx Inc. Nicholas Mattei was supported in part by NSF Awards IIS-RI-2339880, IIS-RI-2007955, IIS-III-2107505, and IIS-RI-2134857. The authors would like to thank Liam Healy for early tinkering and productive discussions and explorations around this topic.
@@ Appendix A Additional Voting Rules
IRV is a deterministic, iterative voting rule that, in each of $m-1$ rounds, eliminates the candidate with the lowest plurality score and removes them from the preference orders of all voters before the next round. When candidates are tied for lowest plurality score we break ties in lexicographically. The rule returns the lottery that assigns all probability to the single candidate that was never eliminated; $U(W(X))$ where $|W(X)|=1$.
@@
If the profile $X$ admits a Condorcet winner $c$, then let $W(X)={c}$. Otherwise, if there is no Condorcet winner, let $W(X)$ be the subset of candidates with maximum Borda score $B(c)$. The probabilistic Black’s rule returns the lottery $U(W(X))$.
@@ Appendix B Rule Preservation
Plurality and Borda are scoring rules, which are necessarily computable from a rank frequency embedding. However neither rule is preserved by the tournament embedding. Plurality is known not to be preserved by the weighted tournament either, but for Borda this remains an open question.
@@
Challenge.
Does the weighted tournament embedding preserve Borda?
@@
Challenge.
Does the weighted tournament embedding preserve IRV?
@@
Appendix C Training Loss for PSCFs
In this Appendix we give the graphs of validation loss for the other rules studied in this paper.
Appendix D Scaling to Fewer Voters
This table gives the results of our scaling experiments to a smaller number of voters.
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.004252 0.148373 0.146005
Borda 0.004678 0.004018 0.075410
Copeland 0.070351 0.051713 0.000698
Schulze 0.080797 0.066160 0.045688
Simpson-Kramer 0.081472 0.061241 0.046492
IRV 0.095856 0.114207 0.088140
Black’s Rule 0.033607 0.033087 0.044673
Table 6: Testing generalization of our models trained to learn PSCFs on uniformly random profiles of 7 candidates and 29 voters on random profiles with $n=9$ voters. Average L1 losses over 1,000 random profiles.
Appendix E Participation-Adjustment
In this section we give the validation losses for learning participation by retraining our models for a given voting rule.

### v2

@@ 7 Conclusions and Future Work
We have shown that not only can we efficiently and accurately learn known PSCFs from preference data, but also that we can fine-tune these rules in order to improve them in ways that, to date, have not be possible through traditional algorithmic design methods. We have highlighted the importance of the choice of embedding on the efficiency of learning and quality of the learned rules. For Participation, we saw that our models trained only on rules did reasonably at satisfying the axiom. After fine tuning, all rules are empirically Participation-proof with minimal loss in rule performance. These adjustment for improving axiomatic properties would be much more difficult, or intractable, to design by hand. Interestingly, we showed none of the tested embeddings retain all the information necessary for IRV, and we see that IRV is indeed the hardest rule to learn across all our testing, reinforcing the importance of embedding/target selection. It remains to be seen whether other embeddings can be designed, of size $m\times m$ or smaller, that outperform the embeddings we took from the social choice literature. Different embeddings may be beneficial in particular for rules whose outcomes are NP-Hard to compute.
@@ Acknowledgments
Matone, Armstrong, and Mattei were supported in part by NSF Awards IIS-RI-2007955, IIS-III-2107505, IIS-RI-2134857, IIS-RI-2339880 and CNS-SCC-2427237 as well as the Harold L. and Heather E. Jurist Center of Excellence for Artificial Intelligence at Tulane University and the Tulane University Center of Excellence for Community-Engaged Artificial Intelligence (CEAI). Ben Abramowitz was supported by the NSF under Grant #2127309 to the Computing Research Association for the CIFellows Project.
Appendix For: DeepVoting: Learning and Fine Tuning Voting Rules with Canonical Embeddings
@@ Appendix A Additional Voting Rules
IRV is a deterministic, iterative voting rule that, in each of m1m-1 rounds, eliminates the candidate with the lowest plurality score and removes them from the preference orders of all voters before the next round. When candidates are tied for lowest plurality score we break ties in lexicographically. The rule returns the lottery that assigns all probability to the single candidate that was never eliminated; $U(W(X))$ where $|W(X)|=1$.
@@
If the profile $X$ admits a Condorcet winner $c$, then let W(X)=cW(X)={c}. Otherwise, if there is no Condorcet winner, let $W(X)$ be the subset of candidates with maximum Borda score $B(c)$. The probabilistic Black’s rule returns the lottery $U(W(X))$.
@@ Appendix B Rule Preservation
Plurality and Borda are scoring rules, which are necessarily computable from a rank frequency embedding. However neither rule is preserved by the tournament embedding. Plurality is known not to be preserved by the weighted tournament either.
$T_{RF}$ $_{WT} TT$$_{T}$
Plurality $\times$ $\times$
Borda $\times$
Copeland $\times$
Schulze $\times$ $\times$
Simpson-Kramer $\times$ $\times$
IRV $\times$ $\times$ $\times$
Black’s Rule $\times$ $\times$
Table 5: PSFC preservation under embedding.
@@
Theorem 10.
The weighted tournament embedding does not preserve IRV.
This follows as a corollary to Halpern, Hossain, and Tucker-Foltz (2024). Thanks to Daniel Halpern for pointing this out!
@@
Appendix C Other Tested Neural Network Setups
We conducted several experiments to explore the efficiency of learning and target larger profile dimensionalities without making significant changes to our models or embeddings. Experimenting with our model’s representational threshold, we found that with higher dimensionalities from our initial baseline of 7 candidates and 29 voters we were able to achieve faster and more consistent divergence with L1Loss and TanH activation.
With 15 candidates and 44 voters, we observe training loss decrease much more consistently than our ReLU baseline with lower variance and better generalization. Similarly, we find that utilization of a distribution-based loss like KLDiv or JensenShannon loss is inefficient when targeting a sparse output vector, and that even with 7 candidates these approaches see a severe reduction in the generalizability of our models.
While we did want to focus on smaller model architectures to more directly rely on our embeddings, we did experiment with more complex models. We tested two alternate networks which we observed to have below-satisfactory performance when compared to our base feed-forward network. The first was a simple network with a depth of 6 and a constant layer width of $n^{2}\times 3$, and the second had the same depth, but a funnel architecture which decomposed the width of each hidden layer from the input to the output layer to minimize information dropoff. Both networks underperformed when compared to our baseline.
Appendix D Scaling Number of Voters
Scaling With Number of Voters
Our embeddings give us the ability to work with profiles with different numbers of voters. Although we trained our models on profiles with 44 voters, we can test with larger and smaller numbers of voters to check generalizability. We test first on profiles with 199 votes (Table 6), and then again with only 13 voters (Table 7).
We see that there is a very mild increase in the loss across most rules and embeddings, though losses remain very similar to their value on profiles with 44 voters (see Table 1). In some cases there is no increase in loss: With the concatenated embedding, $T_{CO}$, both Copeland (with 13 voters) and Black’s rule (199 voters) have the same loss, as does. Loss even decreases when evaluating Black’s rule on profiles with 13 voters and $T_{T}$. As well, on all but 4 rule-embedding pairs, loss is lower using profiles with 13 voters than those with 199 voters. These results highlight the importance of choosing embeddings that fit the rule, corresponding to the learning objective. As a result of our embeddings our learned rules are able to generalize extremely effectively to profiles of varying sizes without the need for additional training.
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.04 0.133 0.142 0.058
Borda 0.165 0.059 0.039 0.057
Copeland 0.095 0.0 0.159 0.001
Schulze 0.08 0.057 0.082 0.057
Simpson-Kramer 0.083 0.057 0.078 0.057
IRV 0.166 0.112 0.166 0.165
Black’s Rule 0.061 0.054 0.166 0.045
Ranked Pairs 0.164 0.059 0.164 0.078
Table 6: L1 Loss on 512 Impartial Culture profiles for networks trained on each embedding 11 candidates and 199 voters.
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.023 0.119 0.131 0.019
Borda 0.166 0.051 0.024 0.047
Copeland 0.095 0.001 0.159 0.0
Schulze 0.083 0.05 0.082 0.049
Simpson-Kramer 0.079 0.049 0.076 0.05
IRV 0.158 0.108 0.158 0.16
Black’s Rule 0.055 0.042 0.166 0.037
Ranked Pairs 0.165 0.055 0.165 0.072
Table 7: L1 Loss on 512 Impartial Culture profiles for networks trained on each embedding 11 candidates and 13 voters.
Appendix E Testing Novel Distributions
This section contains a brief definition of each distribution we tested our networks on and additional results showing the result of testing our networks on each embedding. We include one table for each embedding, showing the effectiveness of the learned embedding-rule pair on data sampled from each test distribution.
Preference Distributions
We train all of our results only on the Impartial Culture which generates profiles with candidates ranked uniformly at random, however we test our networks on each of the following:
  • Impartial Culture (IC) Each unique preference order is equally likely, regardless of which orders any other voters have selected  (Guilbaud 1952).
  • Impartial Anonymous Culture (IAC) Preferences profiles are generated collectively, rather than as individual preference orders. Each multi-set of preference orders (i.e. a preference profile) is equally likely to be generated, making voter identities irrelevant  (Gehrlein and Fishburn 1976).
  • Mallows Preference orders are noisy estimates of some reference ranking $r$, with the amount of noise related to a parameter $\phi$  (Mallows 1957). A value of $\phi=0$ results in all voters have identical preferences while $\phi=1$ results in the Impartial Culture distribution. For each profile we sample $\phi\in[0,1]$ uniformly at random using the Mallow’s distribution described by  Boehmer et al. (2021).
  • Urn All $m!$ preference orders exist in an “urn.” Each voter decides their ranking by sampling a ranking from the urn. Once a ranking is selected, $\alpha!$ copies of it are added to the urn  (Eggenberger and Pólya 1923). For each profile we sample $\alpha$ from a Gamma distribution with shape parameter $k=0.8$ and scale parameter $\theta=1$ as described by Boehmer et al. (2021).
  • Single-Peaked There is some global ordering of alternatives. Each voter has some favourite alternative and prefers all alternatives closer to their favourite over those further away. We sample single-peaked profiles from Walsh’s distribution (Walsh 2015).
  • PrefLib An online repository containing profiles corresponding to real human preferences expressed across many domains  (Mattei and Walsh 2013). We use all complete profiles with strict orders and $m\geq 11$ candidates. For profiles with greater than $11$ candidates we form a profile by selecting a subset of candidates uniformly at random. This results in a test set of 6945 profiles.
Loss and Accuracy
The following tables show L1 Loss and test accuracy on 10,000 test profiles (except in the case of PrefLib) sampled from each of the above distributions. We exclude from our data profiles where any rule results in multiple, tied winners.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.0 0.0 0.032 0.002 0.05 0.021
Borda 0.165 0.165 0.162 0.156 0.179 0.158
Copeland 0.088 0.088 0.109 0.116 0.181 0.112
Schulze 0.071 0.071 0.094 0.106 0.018 0.101
SK 0.071 0.071 0.094 0.106 0.176 0.104
IRV 0.159 0.16 0.163 0.165 0.182 0.165
Black’s Rule 0.046 0.046 0.076 0.079 0.171 0.075
Ranked Pairs 0.161 0.162 0.164 0.165 0.182 0.166
Table 8: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{RF}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.999 1.0 0.851 0.902 0.74 0.883
Borda 0.086 0.085 0.111 0.116 0.0 0.136
Copeland 0.552 0.554 0.408 0.431 0.0 0.379
Schulze 0.719 0.716 0.507 0.498 0.931 0.444
SK 0.72 0.716 0.502 0.489 0.017 0.431
IRV 0.118 0.114 0.097 0.092 0.0 0.092
Black’s Rule 0.767 0.765 0.574 0.596 0.048 0.588
Ranked Pairs 0.102 0.097 0.097 0.091 0.0 0.086
Table 9: Test Accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{RF}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.129 0.131 0.084 0.054 0.09 0.065
Borda 0.048 0.05 0.06 0.007 0.033 0.019
Copeland 0.0 0.0 0.0 0.0 0.0 0.0
Schulze 0.043 0.042 0.025 0.005 0.001 0.01
SK 0.043 0.043 0.026 0.005 0.006 0.01
IRV 0.105 0.105 0.099 0.082 0.086 0.084
Black’s Rule 0.044 0.045 0.028 0.005 0.008 0.011
Ranked Pairs 0.051 0.051 0.036 0.017 0.001 0.02
Table 10: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{T}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.355 0.35 0.599 0.633 0.493 0.642
Borda 0.826 0.82 0.704 0.902 0.856 0.895
Copeland 1.0 0.999 1.0 1.0 1.0 0.981
Schulze 0.903 0.908 0.93 0.967 1.0 0.934
SK 0.902 0.906 0.931 0.975 0.999 0.933
IRV 0.446 0.448 0.489 0.531 0.555 0.537
Black’s Rule 0.866 0.86 0.914 0.96 0.994 0.94
Ranked Pairs 0.814 0.818 0.863 0.897 1.0 0.888
Table 11: Test accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{T}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.136 0.137 0.122 0.109 0.094 0.113
Borda 0.018 0.019 0.048 0.078 0.018 0.071
Copeland 0.163 0.162 0.157 0.151 0.182 0.152
Schulze 0.067 0.067 0.118 0.138 0.137 0.133
SK 0.062 0.061 0.114 0.136 0.17 0.13
IRV 0.159 0.16 0.163 0.165 0.182 0.165
Black’s Rule 0.165 0.164 0.166 0.166 0.182 0.166
Ranked Pairs 0.161 0.162 0.164 0.165 0.182 0.166
Table 12: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{WT}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.289 0.29 0.355 0.388 0.537 0.379
Borda 0.934 0.935 0.737 0.666 0.935 0.611
Copeland 0.087 0.092 0.138 0.148 0.0 0.166
Schulze 0.735 0.744 0.355 0.338 0.287 0.281
SK 0.79 0.794 0.378 0.355 0.013 0.27
IRV 0.118 0.114 0.098 0.092 0.0 0.092
Black’s Rule 0.088 0.092 0.086 0.081 0.0 0.073
Ranked Pairs 0.102 0.097 0.097 0.091 0.0 0.085
Table 13: Test Accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{WT}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.001 0.001 0.018 0.012 0.031 0.014
Borda 0.028 0.03 0.077 0.093 0.135 0.109
Copeland 0.0 0.0 0.001 0.001 0.0 0.002
Schulze 0.017 0.016 0.029 0.031 0.018 0.042
SK 0.017 0.016 0.028 0.027 0.02 0.039
IRV 0.163 0.162 0.163 0.164 0.182 0.162
Black’s Rule 0.022 0.023 0.066 0.085 0.018 0.103
Ranked Pairs 0.06 0.06 0.061 0.06 0.001 0.063
Table 14: L1 Loss across preference distributions of networks trained using Impartial Culture preferences and the $T_{CO}$ embedding.
Target Rule IC IAC Urn Mall. SP PrefLib
Plurality 0.997 0.997 0.9 0.942 0.847 0.911
Borda 0.844 0.834 0.577 0.491 0.263 0.4
Copeland 1.0 1.0 0.999 0.998 1.0 0.977
Schulze 0.914 0.919 0.881 0.884 0.946 0.821
SK 0.913 0.917 0.888 0.914 0.894 0.829
IRV 0.106 0.109 0.101 0.1 0.0 0.111
Black’s Rule 0.88 0.873 0.638 0.534 0.9 0.433
Ranked Pairs 0.67 0.672 0.667 0.67 0.996 0.653
Table 15: Test accuracy across preference distributions of networks trained using Impartial Culture preferences and the $T_{CO}$ embedding.
Appendix F Participation Fine-Tuning
In this section we show the test loss for each of our rule-embedding pairs before and after fine-tuning (Table 16, Table 17, Table 18). Loss is calculated on a set of 1056 profiles with 44 voters. Across $T_{RF}$ and $T_{T}$ embeddings most rules experience only a minor increase in Rule Loss in exchange for significant decrease in Participation Loss during fine-tuning. Curiously, the $T_{WT}$ embedding is an exception to this; our fine-tuning process appears to optimize heavily for minimizing Participation Loss while greatly increasing Rule Loss. This suggests to us that the Weighted Tournament embedding provide information particularly well-suited to Participation Loss while being too complex for learning many rules. We also plot training loss during fine-tuning for each rule-embedding pair. Each plot shows 40 epochs of fine-tuning with a separate series for Rule Loss (L1 Loss), Participation Loss, and the sum of both loss terms. While Rule Loss occasionally increases a moderate amount during fine-tuning we see that combined loss consistently drops and in almost all cases the increase to Rule Loss is quite mild.
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.0 0.247 0.095 0.005
Borda 0.165 0.0 0.165 0.0
Copeland 0.084 0.274 0.168 0.0
Schulze 0.064 0.352 0.076 0.069
Simpson-Kramer 0.068 0.313 0.074 0.097
IRV 0.158 0.0 0.158 0.0
Black’s Rule 0.037 0.291 0.161 0.001
Ranked Pairs 0.159 0.0 0.159 0.0
Table 16: Loss of our $T_{RF}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.098 0.528 0.113 0.16
Borda 0.034 0.352 0.042 0.226
Copeland 0.0 0.261 0.023 0.144
Schulze 0.017 0.361 0.036 0.132
Simpson-Kramer 0.02 0.358 0.037 0.139
IRV 0.099 0.27 0.102 0.184
Black’s Rule 0.027 0.339 0.037 0.211
Ranked Pairs 0.044 0.305 0.057 0.231
Table 17: Loss of our $T_{T}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.102 0.555 0.164 0.013
Borda 0.009 0.296 0.162 0.0
Copeland 0.163 0.0 0.162 0.0
Schulze 0.065 0.276 0.165 0.0
Simpson-Kramer 0.057 0.299 0.163 0.0
IRV 0.158 0.0 0.158 0.0
Black’s Rule 0.165 0.0 0.165 0.0
Ranked Pairs 0.159 0.0 0.159 0.0
Table 18: Loss of our $T_{WT}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
Figure 4: Training Loss during fine-tuning for each rule using the $T_{RF}$ embedding.
Figure 5: Training Loss during fine-tuning for each rule using the $T_{T}$ embedding.
Figure 6: Training Loss during fine-tuning for each rule using the $T_{WT}$ embedding.
Figure 7: Training Loss during fine-tuning for each rule using the $T_{CO}$ embedding.